OpenStudy (anonymous):

How much would $200 invested at 4% be worth after 8 years

OpenStudy (amistre64):

depends on how the interest is compounded ...

OpenStudy (amistre64):

if its simple interest, then A = P(1+rt), otherwise it gets messy

OpenStudy (amistre64):

i was kinda hoping that the asker could fill in some of the parts, and participate in the solution process.

OpenStudy (texaschic101):

I am sorry....I will delete my comments

OpenStudy (wolf1728):

Let's assume 4% compounded annually Amount Earned = Principal * (1+rate)^years Amount Earned = $200 * (1.04)^8 Amount Earned = $200 * 1.3685690504 Amount Earned = $273.7138100811 Amount Earned = $273.71 Compound Interest Calculator to check your work: http://1728.org/compint.htm